OpenStudy (anonymous):

Global sea level is rising about 3 mm/yr. How many $km^3$ of continental ice would have to melt in one century in order to supply the fresh water needed to account for this rate of sea level rise. Where on Earth would most of this ice-melt come from?

OpenStudy (anonymous):

3mm/year=300mm/century \[\frac{ 300mm }{ century } \times \frac{ 1km }{ 1,000,000mm} \times \frac{ 1km }{ 1,000,000mm } \times \frac{ 1km }{ 1,000,000 }\]

OpenStudy (anonymous):

It would come from glacier at the North and South Pole or upper and low latitudes